OpenStudy (agl202):

Medal Fan for best answer!!! Line AB contains points A (-6, 3) and B (2, -5). Line AB has a slope that is... zero undefined positive negative

OpenStudy (johnp12):

0

OpenStudy (johnp12):

i thimk

OpenStudy (anonymous):

Find the slope. \(m = \dfrac{y_2-y_1}{x_2-x_1}\) \(m = \dfrac{-5-3}{2+6}\) Subtract & add

OpenStudy (anonymous):

@johnp12 Don't give answers, it's against the Code of Conduct.

OpenStudy (anonymous):

Plus your answer is wrong.

OpenStudy (anonymous):

What's -5 - 3? @Agl202

OpenStudy (agl202):

Is the answer D. then?

OpenStudy (agl202):

@iGreen. ?

OpenStudy (holly00d1248):

@iGreen. is the slope 1

OpenStudy (holly00d1248):

@Agl202 your not done yet

OpenStudy (anonymous):

Yep, you got it.

OpenStudy (anonymous):

You get -8/8 and that gives you -1.

OpenStudy (agl202):

Yep, thats what I got! Thx... for ur help!!! :)

OpenStudy (anonymous):

No problem!